What is the bulk modulus of oxygen if 38.7 g of oxygen occupies 25.5 L and the speed of sound in the oxygen is 342 m/s?
Answer #1

By applying density formula,

p = m/V

= 0.0387 kg/0.0255 m^3 ( 1 L = 0.001 m^3)

= 1.518 kg/m^3

We know that,

The bulk modulus is related to the density of the gas and the speed of sound in the gas

K = p*c^2

= 1.518 kg/m^3 * (342 m/s)^2

= 1.775 * 10^5 Pa

= 177 kPa
Answer